OpenStudy (thadds2003):

What is the ratio 15 : 75 in lowest terms? A. 1 : 5 B. 2 : 5 C. 1 : 4 D. 3 : 5

OpenStudy (thadds2003):

im thinking c.

OpenStudy (anonymous):

yes you are correct!

OpenStudy (thadds2003):

confirmation! @sleepyjess

OpenStudy (baru):

what is 15*4 =?

OpenStudy (baru):

15*1=15 15*5=75

OpenStudy (thadds2003):

umm.... 60!

OpenStudy (baru):

yea 15*4=60 so why u think its option c?